OpenStudy (anonymous):

Which literal equations are equivalent to F = ma? Choose all answers that are correct. A. B. C. D.

OpenStudy (agl202):

So what is the options for the answers???

OpenStudy (agl202):

@vmeza0131

OpenStudy (agl202):

m= F/A and A= F/m can be derived from the force equation (where the sum of all forces acting on an object or system is equal to its mass times its net acceleration).

OpenStudy (agl202):

Did this help???

OpenStudy (anonymous):

sorry I haven't logged in in over 24 hours but I got it maybe later I will need help @Agl202

OpenStudy (agl202):

Oh okay